# MORE: Multi-mOdal REtrieval Augmented Generative Commonsense Reasoning

### v1

@@ Abstract
Wanqing Cui, Keping Bi, Jiafeng Guo, Xueqi Cheng CAS Key Lab of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China {cuiwanqing18z, bikeping, guojiafeng, cxq}@ict.ac.cn
@@ 3.2.1 Retrieval Results for Augmentation
Specifically, after we preprocessed the retrieved results by removing duplicate images and noisy text, we collected a total of 500,100 images and 787,970 text snippets. On average, each concept set has 14 images and 23 passages, which means that and can be at most 14 and 23 respectively. See the Appendix B for more details.
@@ 4 Experiments Settings
| Model | Bleu | METEOR | ROUGE | CIDEr | SPICE |
| gpt-3.5-turbo (0 shot) | 21.44 | 28.93 | 48.80 | 13.03 | 26.69 |
| gpt-3.5-turbo (3 shot) | 28.91 | 31.14 | 53.25 | 15.92 | 28.89 |
@@
| MORE-T5 (text) | 29.87* | 30.15* | 55.20* | 15.70* | 31.57* |
| MORE-T5 (image) | 29.98* | 30.21* | 55.07* | 15.92* | 31.63* |
| MORE-T5 (multi-modal) | 30.27* | 30.28* | 55.18* | 16.02* | 31.94* |
@@
| MORE-T5 (text) | 32.03* | 31.05* | 56.14 | 16.37 | 32.00 |
| MORE-T5 (image) | 32.37* | 31.31* | 56.60* | 16.67* | 32.31* |
| MORE-T5 (multi-modal) | 32.29* | 30.90* | 56.62* | 16.63* | 32.34* |
@@
| MORE-OPT (multi-modal) | 32.78* | 32.15 | 57.07* | 17.03* | 32.94* |
Model Bleu CIDEr SPICE GPT-4 28.53 16.41 29.05 GPT-4 (1i1t) 19.86 12.43 26.2 GPT-3.5 (0 shot) 22.97 13.93 27.25 GPT-3.5 (3 shot) 28.35 16.14 29.13 MORE (1i1t) 31.81* 17.08* 31.81* MORE (3i3t) 32.53* 17.30* 32.81*
4.1 Dataset
4.2 Methods for Comparisons
Text-based/Multi-modal LMs: For text-based LMs, we employ T5 as well as T5 Raffel et al. (2019) to represent small pre-trained LMs, and OPT Zhang et al. (2022) to represent the LLMs. We also query the close source model gpt-3.5-turbo OpenAI (2023a) through API with the prompt "Use the given words to make a short sentence that is consistent with commonsense. Words: {…}". For Multi-modal LMs (MLMs) we compare with BLIP2-OPT Li et al. (2023), an open source model, and gpt-4-vision-preview OpenAI (2023b), a close source model. Since MLMs can accept images and text as input, we directly input the retrieved items into MLMs. All the above open source models are based on huggingface222https://github.com/huggingface and are under Apache License 2.0.
Retrieval Augmented Generation Baselines: We consider two types of textual retrieval augmented models. One is Prepend Mallen et al. (2023); Si et al. (2023); Ram et al. (2023), which prepends the top-k text results to the concepts as input. The other one is FiD Izacard and Grave (2021), which concatenates each retrieved passage with the concept words separately to encode in parallel for better handling of long text. For the visual retrieval augmented model, we compare with VisCTG Feng et al. (2022), which generates a caption for each image with an image captioning model Luo et al. (2018) and pretends the captions to the input for augmentation. All the above models use T5 as the backbone and are tuned with prompt-tuning.
MORE with Various Backbones: We test MORE with various backbones to explore whether it can be effectively used in different model architectures. Specifically, T5 and T5 represent small LMs and are encoder-decoder architecture. BLIP2-OPT represent MLMs. It should be noted that BLIP2-OPT is equivalent to OPT when not receiving image input. Therefore it can also be regarded as a variant based on OPT, which represents LLMs and is decoder-only architecture.
@@ 4.3 Evaluation Metrics
To assess the generation performance, we use standard metrics: BLEU Papineni et al. (2002) quantifying the overlap between predictions and references based on n-gram precision and ROUGE Lin (2004) measuring the n-gram recall. METEOR Banerjee and Lavie (2005) is an improved version of BLEU and considers both exact word matches and semantic similarities. CIDEr Vedantam et al. (2014) focuses on capturing sentence semantic similarity. SPICE Anderson et al. (2016) quantifies the semantic propositional content of generations by leveraging scene graphs and this score is most consistent with human evaluation. We use the entire test dataset to obtain the main results and randomly sample 500 data in the test set to compare with GPT-4 for the sake of limit budget.
@@ 4.4 Implementation Details
| Model | Bleu | CIDEr | SPICE |
@@
| Model | Bleu | CIDEr | SPICE |
| T5 (32 pl) | 28.93 | 15.36 | 30.95 |
| T5 (64 pl) | 28.74 | 14.84 | 30.9 |
| MORE (rand-RA) | 29.33 | 15.54 | 30.73 |
| MORE | 30.27 | 16.02 | 31.94 |
5 Results and Analysis
5.1 Overall Results
As shown in Table 1, after incorporating text and images to LMs, our method can boost the generation performance significantly based on various backbones. Comparing images and text, we find that images are better in improving commonsense ability, and incorporating both of them yields even better performance.
As shown in Table 2, Although based on a smaller model, MORE can achieve better results than GPT-3.5 and GPT-4. This fully illustrates the effectiveness of our method. Considering that GPT-4 is a multi-modal language model, we also test its performance when retrieval augmented items are provided. However, GPT-4 cannot effectively utilize the retrieved inputs, leading to deteriorated performance. Retrieval augmentation methods for the GPT-4 model are worth exploring in the future.
@@ Appendix A Test Result Using Captions
| Model | Bleu | CIDEr | SPICE |
@@ Appendix B Retrieved Inputs Crawling and Preprocessing
For the image, we use the template ‘a photo of {…}’ and crawl the first 20 image results returned by the search engine. We further removed duplicate images based on the dHash algorithm. For text, we crawl the results from the first two pages. Considering that the full document associated with each result may be very long, the search engine has provided a concise text summary of the webpage aligning with the search keywords, we only keep the title and description in the snapshot. We also removed the URL and non-English parts.

### v2

@@ Abstract
Wanqing Cui, Keping Bi††thanks: Corresponding author., Jiafeng Guo, Xueqi Cheng CAS Key Lab of Network Data Science and Technology, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China {cuiwanqing18z, bikeping, guojiafeng, cxq}@ict.ac.cn
@@ 3.2.1 Retrieval Results for Augmentation
Specifically, after we preprocessed the retrieved results by removing duplicate images and noisy text, we collected a total of 500,100 images and 787,970 text snippets. On average, each concept set has 14 images and 23 passages, which means that and can be at most 14 and 23 respectively. We retain the order returned by the browser without any re-ranking. See the Appendix B for more details.
@@ 4 Experiments Settings
4.1 Dataset
4.2 Methods for Comparisons
Text-based/Multi-modal LMs: For text-based LMs, we employ T5 as well as T5 Raffel et al. (2019) to represent small pre-trained LMs, and OPT2.7b Zhang et al. (2022) to represent the LLMs. We also query the close source model gpt-3.5-turbo-1106 OpenAI (2023a) through API with the prompt "Use the given words to make a short sentence that is consistent with commonsense. Words: {…}". For Multi-modal LMs (MLMs) we compare with BLIP2-OPT2.7b Li et al. (2023), an open source model, and gpt-4-1106-vision-preview OpenAI (2023b), a close source model. Since MLMs can accept images and text as input, we directly input the retrieved items into MLMs. All the above open source models are based on huggingface222https://github.com/huggingface and are under Apache License 2.0.
Retrieval Augmented Generation Baselines: We consider two types of textual retrieval augmented models. One is Prepend Mallen et al. (2023); Si et al. (2023); Ram et al. (2023), which prepends the top-k text results to the concepts as input. The other one is FiD Izacard and Grave (2021), which concatenates each retrieved passage with the concept words separately to encode in parallel for better handling of long text. For the visual retrieval augmented model, we compare with VisCTG Feng et al. (2022), which generates a caption for each image with an image captioning model Luo et al. (2018) and prepends the captions to the input for augmentation. All the above models use T5 as the backbone and are tuned with prompt-tuning.
MORE with Various Backbones: We test MORE with various backbones to explore whether it can be effectively used in different model architectures. Specifically, T5 and T5 represent small LMs and are encoder-decoder architecture. BLIP2-OPT2.7b represent MLMs. It should be noted that BLIP2-OPT2.7b is equivalent to OPT2.7b when not receiving image input. Therefore it can also be regarded as a variant based on OPT2.7b, which represents LLMs and is decoder-only architecture.
| Model | Bleu4 | METEOR | ROUGEL | CIDEr | SPICE |
| GPT-3.5 (0-shot) | 21.44 | 28.93 | 48.80 | 13.03 | 26.69 |
| GPT-3.5 (3-shot) | 28.91 | 31.14 | 53.25 | 15.92 | 28.89 |
@@ 4.2 Methods for Comparisons
| MORE (text) | 29.87* | 30.15* | 55.20* | 15.79* | 31.57* |
| MORE (image) | 29.98* | 30.21* | 55.07* | 15.92* | 31.63* |
| MORE (multi-modal) | 30.27* | 30.28* | 55.18* | 16.02* | 31.94* |
@@
| MORE (text) | 32.03* | 31.05* | 56.14 | 16.37 | 32.00 |
| MORE (image) | 32.37* | 31.31* | 56.60* | 16.67* | 32.31* |
| MORE (multi-modal) | 32.29* | 30.90* | 56.62* | 16.63* | 32.34* |
@@
| MORE (multi-modal) | 32.78*† | 32.15† | 57.07*† | 17.03*† | 32.94*† |
@@ 4.3 Evaluation Metrics
To assess the generation performance, we use standard metrics: BLEU Papineni et al. (2002) quantifying the overlap between predictions and references based on n-gram precision and ROUGE Lin (2004) measuring the n-gram recall. METEOR Banerjee and Lavie (2005) is an improved version of BLEU and considers both exact word matches and semantic similarities. CIDEr Vedantam et al. (2014) focuses on capturing sentence semantic similarity. SPICE Anderson et al. (2016) quantifies the semantic propositional content of generations by leveraging scene graphs. Please notice that SPICE aligns closely with human evaluation and should be treated as the primary metric. We also incorporate sentence similarity metrics (Sent-Sim) with SimCSE Gao et al. (2021) to measure semantic similarity. We use the entire test dataset to obtain the main results and randomly sample 500 data in the test set to compare with GPT-4 for the sake of a limited budget.
@@ 4.4 Implementation Details
T5 MORE T5 MORE OPT BLIP2 MORE GPT-3.5 (3-shot) Avg 71.51 71.59 72.12 72.23 71.66 71.83 72.53 70.00 Max 82.32 83.31 83.9 84.05 83.15 83.8 84.15 80.08
5 Results and Analysis
5.1 Overall Results
As shown in Table 1 and Table 2, after incorporating text and images to LMs, our method can boost the generation performance significantly based on various backbones. Comparing images and text, we find that images are better in improving commonsense ability, and incorporating both of them yields even better performance.
As shown in Table 3, Although based on a smaller model, MORE can achieve better results than GPT-3.5 and GPT-4. This fully illustrates the effectiveness of our method. Considering that GPT-4 is a multi-modal language model, we also test its performance when retrieval augmented items are provided. However, GPT-4 cannot effectively utilize the retrieved inputs, leading to deteriorated performance. Retrieval augmentation methods for the GPT-4 model are worth exploring in the future. We also tested the model with a specified length limit to avoid the tendency of LLMs to generate longer sentences. The length of each sentence is the average length of the golden references, so the results can be regarded as an upper bound. According to the most critical SPICE metric, length constraints do not lead to better results. Further analysis revealed that length constraints result in the concept coverage decrease, indicating that LLMs face challenges in organizing concepts with simple short sentences.
Model Bleu4 CIDEr SPICE GPT-4 (0-shot) 28.53 16.52 30.53 GPT-4 (0-shot & lc) 27.87 16.89 29.11 GPT-4 (3-shot) 30.00 16.41 29.05 GPT-4 (0-shot & 1i1t) 19.86 12.43 26.20 GPT-3.5 (0-shot) 22.97 13.93 27.25 GPT-3.5 (0-shot & lc) 25.54 15.62 26.75 GPT-3.5 (3-shot) 28.35 16.14 29.13 MORE (1i1t) 31.81*† 17.08*† 31.81*† MORE (3i3t) 32.53*† 17.30*† 32.81*†
| Model | Bleu4 | CIDEr | SPICE |
@@ 5.3 Analysis
| Model | Bleu4 | CIDEr | SPICE |
| T5 (32 pl) | 28.93 | 15.36 | 30.95 |
| T5 (64 pl) | 28.74 | 14.84 | 30.9 |
| MORE (rand-RA) | 29.33 | 15.54 | 30.73 |
| MORE | 30.27 | 16.02 | 31.94 |
@@ 7 Limitations
Acknowledgement
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62302486, the Innovation Project of ICT CAS under Grants No. E361140, the CAS Special Research Assistant Funding Project, the Lenovo-CAS Joint Lab Youth Scientist Project, the project under Grants No. JCKY2022130C039, and the Strategic Priority Research Program of the CAS under Grants No. XDB0680102.
@@ Appendix A Test Result Using Captions
| Model | Bleu4 | CIDEr | SPICE |
@@ Appendix B Retrieved Inputs Crawling and Preprocessing
We concatenate all concepts in a concept set to form a query. For the image, we use the template ‘a photo of {…}’ (e.g. a photo of decorate, music, background, and tree) and crawl the first 20 image results returned by the search engine. We further removed duplicate images based on the dHash algorithm. For text, we directly use the concatenation of the concept set as the query. We crawl the text results from the first two pages. Considering that the full document associated with each result may be very long, the search engine has provided a concise text summary of the webpage aligning with the search keywords, we only keep the title and description in the snapshot. We also removed the URL and non-English parts.